# MVGT: A Multi-view Graph Transformer Based on Spatial Relations for EEG Emotion Recognition

### v2

@@ IV-C2 Geometric Structure Encoding
In the real world, the human reasoning process considers not only the semantic relationships between objects but also their spatial relations. EEG channels have a 3D structure, and the functional connectivity between these channels lack precise definitions. Therefore, we represent the relationships between EEG channels as a fully connected directed graph structure. The Euclidean distance between channels is calculated using their coordinates to learn the spatial correlations between nodes. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [43, 30]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
where $i$ is the node index, $k$ is the index of basis function and $\bm{W}_{\mathcal{Z}}\in\mathbb{R}^{K\times d}$ is a learnable projection function. $\text{Projection}:\mathbb{R}^{n\times n\times K}\mapsto\mathbb{R}^{n\times n\times M}$ is a nonlinear transformation, where $M$ is the number of attention heads. We incorporate this encoding as a bias term into the softmax attention.
Our proposed spatial encoding matrix is directed, which is inconsistent with the assumption of a symmetric adjacency matrix [16, 13]. Using directed connections provides the model with greater expressive power because the correlation between node pairs $\left(i,j\right)$ and $\left(j,i\right)$ may differ. Since we assume nodes are fully connected, we avoid specific assumptions about inter-channel correlations and learn the functional correlations between nodes through encoding. Let $l$ denote the model depth, and $i$ denote the index of multi-head attention. Therefore, the brain functional encoding of the $s$-th sample can be represented as:
$\bm{A}^{l,i}=\text{Softmax}\left(\frac{\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{Q}}^{l,i}\right)\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{K}}^{l,i}\right)^{T}}{\sqrt{d_{h}^{l,i}}}+\bm{B^{\prime}}^{i}\right),$ (9)
where the projections are learnable parameters $\bm{W}_{\mathcal{Q}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$ and $\bm{W}_{\mathcal{K}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$, and $d_{h}^{l,i}$ is the second dimension of $\bm{W}_{\mathcal{K}}^{l,i}$. The scalar $d_{h}$ represents the number of the hidden dimension. This encoding method integrates temporal, frequency, and spatial domain features into the model, enhancing its expressive power. We compute the attention scores between nodes using embedded vectors, representing the semantic correlations between different nodes from multiple perspectives. Finally, the attention scores are added to the spatial geometric encoding to obtain the correlations between channels.
@@ IV-D Implementation Details of MVGT
In this section, we describe the overall architecture of the model, including spatial encoding and the Transformer encoder, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [44] to normalize the input features between 0 and 1. Subsequently, we perform geometric and regional structure encoding to obtain multi-domain embeddings.
$\bm{X^{\prime}}_{s}$ $=\text{GraphNorm}(\bm{X}_{s}),$ (10)
$\bm{H}^{0}_{s}$ $=\text{SpatialEncoding}+\text{Proj}(\bm{X^{\prime}}_{s}),$ (11)
We employ a Pre-LN Transformer structure, applying layer normalization (LN) before the multi-head attention (MHA) and the FFN. Recent study suggests that the Pre-LN structure yields more stable gradients and is more favorable for optimizer, enabling faster convergence [45] compared to Post-LN. Additionally, we utilize dropout to mitigate overfitting. This process is represented as follows:
$\bm{H^{\prime}}^{l}_{s}$ $=\text{MHA}(\text{LN}(\bm{H}^{l-1}_{s}))+\bm{H}^{l-1}_{s},$ (12)
$\bm{H}^{l}_{s}$ $=\text{FFN}(\text{LN}(\bm{H^{\prime}}^{l}_{s}))+\bm{H^{\prime}}^{l}_{s},$ (13)
Inspired by [30, 46], we feed the outputs recursively into the same modules, denoted as recycling in Fig. 1. The iterative refinement progressively refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
@@ V-A Datasets
The datasets consist of EEG signals recorded from subjects while they watched emotion-eliciting videos.
SEED [5] dataset comprises data from 15 subjects who participated in three sessions, each separated by at least one week. Each sessions consists of 15 trials capturing emotional labels, with the emotion labels being positive, negative, and neutral.
SEED-IV [47] dataset is constituted by EEG signals from 15 subjects across three separate sessions conducted at different times, using the same device as the SEED dataset. This dataset encompasses four emotion labels: neutral, sad, fear, and happy. In each session, each subject underwent 24 trials.
@@ V-B Settings
To prevent potential data leakage that could arise from segment-wise shuffling, we split the training and test sets at the trial level. Following the settings of previous studies [5, 10, 12, 13, 14, 16, 47, 31], we use pre-computed differential entropy (DE) features for the recognition task. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we use the first 16 trials as the training set and the last 8 trials as the test set. The DE features for SEED-IV are calculated using 4s windows. The performance of our model is assessed using data from all three sessions.
For input data, we use overlapping sliding windows of size $T$ along the time axis to extract sample fragments, with $T$ being set to 5. During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. The iterative refinement process is performed three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [48] is employed as the optimizer with a weight decay rate of 0.1.
V-C Results Analysis
We compare the classification results based on the SEED and SEED-IV datasets with recent state-of-the-art models,
Table I: The classification accuracies (mean/std) on SEED and SEED-IV
@@ V-C Results Analysis
as shown in Table 1. It is evident that our proposed model significantly outperforms the baseline models under the same experimental settings. In the experiments on the SEED dataset, the model adopting the FRONTAL scheme achieved the best performance, with a classification accuracy of 96.45%. The LOBE scheme also achieved a slightly superior accuracy of 95.36%, compared to other models. For the SEED-IV dataset, the classification accuracy under the GENERAL scheme was 93.57%, reaching the best performance compared to baseline models. The MVGT model also achieved commendable results under other division schemes. Overall, our model achieved the best recognition accuracy compared to the baselines. The results suggest that selecting the specific division scheme relevant to the emotion task could enhance the expressive power of MVGT.
Figure 3: Confusion matrices of MVGT. (a) Confusion matrix of MVGT-F on the SEED datasets. (b) Confusion matrix of MVGT-G on the SEED-IV datasets. Each row of the matrix represents the ture labels while each column serves as the predicted labels.
Fig. 3 illustrates the confusion matrices of MVGT-F on the SEED and MVGT-G on the SEED-IV, respectively. The values represent the recognition accuracy of the model for different emotion classes. For the SEED dataset, our model achieved the highest accuracy in recognizing positive emotions (98.12%), followed by neutral emotions (96.38%), with negative emotions being slightly lower (94.73%). Only 0.33% of positive emotion samples were misclassified as negative, while only 0.76% of negative emotion samples were recognized as positive, indicating the model’s effectiveness in distinguishing valence changes. For the SEED-IV dataset, our model performed best in recognizing neutral emotions, with an accuracy of 95.90%, while its performance on happy emotions was slightly lower than the other three emotions, with an accuracy of 90.76%. This could be attributed to the GENERAL scheme setting, making the model more sensitive to balanced emotions.
Our model achieved state-of-the-art performance on both the SEED and SEED-IV datasets, primarily due to our comprehensive consideration of frequency, temporal, and spatial geometric information, combined with prior knowledge from neuroscience. The incorporation of relevant brain region schemes into the model significantly contributed to its success.
@@ V-D Ablation Study
To validate the effectiveness of spatial encodings, we conducted ablation experiments on the SEED and SEED-IV datasets, as presented in Table II. By removing both types of spatial encoding, we repeated the aforementioned experiments under the same experimental settings. On the SEED dataset, the model achieved an accuracy of 93.79% with a standard deviation of 7.15%. Compared to MVGT-F, the accuracy decreased by 2.66% and the standard deviation increased by 2.75% after removing spatial encodings. For the SEED-IV dataset, the accuracy dropped by 4.08%, resulting in 89.49%, with the standard deviation rising by 1.80% to 10.40%, when compared to MVGT-G. The experiments demonstrate that incorporating spatial structure information benefits the model performance in emotion recognition tasks. Under experimental settings that consider only geometric structure or brain region structure, the model’s classification accuracy improved over the plain model without any spatial encoding. Evidently, when considering both types of spatial structures simultaneously, the model performance significantly surpassed that of the plain model and models using only single spatial information. This indicates the effectiveness of our proposed spatial encodings and confirms that the expressive power of the Graph Transformer relies on the spatial structure and positional encoding.
Figure 4: The learned inter-channel relationships from the SEED by the MVGT-F and from the SEED-IV by the MVGT-G are illustrated. The figures show the last iteration in the iterative refinement, highlighting the top 10 channel pairs with the highest weights after softmax processing (darker colors indicate higher weights). Channels of the same brain region are represented in the same color. Rows correspond to attention heads, while columns represent the layers of the MHA mechanism.
@@ V-E Visualization of Inter-channel relations
To better illustrate the correlations between channels, we visualized the inter-channel relations of MVGT-F on the SEED and MVGT-G on the SEED-IV. Given that the inter-channel relations might vary among different subjects, we calculated the average weights across all subjects. We focused on the last iteration of iterative refinement and selected the 10 strongest connections of channel pairs. Fig. 4 shows the visualization results, where the rows represent the attention heads and the columns represent the layers of the MHA.
The parameters based on the SEED dataset indicate that emotion patterns are reflected in the activities of multiple brain regions. In the first layer of MVGT-F, the channels in the left frontal region had higher participation in the first attention head, while the channels in the right frontal region were more involved in the second head, potentially corresponding to positive and negative emotion patterns [9], respectively. In the second layer of the model, the parietal and occipital regions showed higher involvement, which aligns with the findings on emotion patterns in [49]. As the model depth increases, the symmetrical connections in the lateral temporal regions of both hemispheres are enhanced, consistent with previous research by [5, 13, 16]. For the SEED-IV dataset, the connections in the frontal, parietal, and occipital regions are the most active, consistent with the findings of [13]. In the first attention head of MVGT-G, the strongest correlation was between O1 and PO3, followed by P4 and P2. Other connections were mainly distributed in the temporal and frontal regions. In the second head, the channel pairs (O1, PO5), (CB1, PO7), and (PO5, PO7) contributed the most to emotion recognition. Additionally, the connection between AF3 and FP1 provided important information for emotion processing, which aligns with the conclusions of [13, 16].
Overall, our model does not focus solely on the local information of a single brain region but comprehensively considers intra-regional and inter-regional information. This confirms that emotional states result from interactions among widely distributed functional networks in the brain, as discussed by [50].
@@ VI Conclusions
In this paper, we propose a multi-view Graph Transformer based on spatial relations for EEG emotion recognition. This model integrates information from multiple perspectives, including temporal, frequency and spatial domains. We incorporate spatial geometric encoding and brain region encoding to enhance the Graph Transformer’s ability to perceive spatial structures. Additionally, the model adaptively learns inter-channel relationships through an attention mechanism and the encoding of channel geometry. Extensive experiments on public emotion recognition datasets demonstrate that our proposed model outperforms other competitive baseline models.
Furthermore, analysis of channel correlations indicates that emotional activities in the brain are not confined to a single local region but result from the coordinated action of multiple brain areas. The frontal, parietal, occipital, and lateral temporal lobes all contribute to the emotion recognition tasks in varying degrees.

### v3

@@ IV-C2 Geometric Structure Encoding
In the real world, the human reasoning process considers not only the semantic relationships between objects but also their spatial relations. EEG channels have a 3D structure, and the functional connectivity between these channels lacks precise definitions. Therefore, we represent the relationships between EEG channels as a fully connected directed graph. Firstly, let $\phi(i,j)$ represent the Euclidean distance between node $i$ and node $j$, and encode $\phi(i,j)$ using a set of Gaussian basis functions [43, 30]. Let $\bm{b}_{k}\in\mathbb{R}^{n\times n}$ denote one of the Gaussian basis functions. The element $(i,j)$ of this function can be expressed as:
@@
where $i$ is the node index, $k$ is the index of basis function and $\bm{W}_{\mathcal{Z}}\in\mathbb{R}^{K\times d}$ is a learnable projection matrix. $\text{Projection}:\mathbb{R}^{n\times n\times K}\mapsto\mathbb{R}^{n\times n\times M}$ is a nonlinear transformation, where $M$ is the number of attention heads. We incorporate this encoding as a bias term into the softmax attention.
Our proposed spatial encoding matrix is directed, which is inconsistent with the assumption of a symmetric adjacency matrix [16, 13]. Using directed connections provides the model with greater expressive power because the correlation between node pairs $\left(i,j\right)$ and $\left(j,i\right)$ may differ. Since we assume nodes are fully connected, we avoid specific assumptions about inter-channel correlations and learn the functional correlations between nodes through the encodings. Letting $l$ denote the model depth, and $i$ denote the index of multi-head attention, the functional brain connectivity of the $s$-th sample can be represented as:
$\bm{A}^{l,i}_{s}=\text{Softmax}\left(\frac{\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{Q}}^{l,i}\right)\left(\bm{H}^{l}_{s}\bm{W}_{\mathcal{K}}^{l,i}\right)^{\top}}{\sqrt{d_{h}^{l,i}}}+\bm{B}^{\prime i}\right),$ (9)
where the projections are learnable parameters $\bm{W}_{\mathcal{Q}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$ and $\bm{W}_{\mathcal{K}}^{l,i}\in\mathbb{R}^{d\times d_{h}}$. The scalar $d_{h}^{l,i}$ is the second dimension of $\bm{W}_{\mathcal{K}}^{l,i}$. This encoding method integrates temporal, frequency, and spatial domain features into the model, enhancing its expressive power. We compute the attention scores between nodes using embedded vectors, representing the semantic correlations between different nodes from multiple perspectives. Finally, the attention scores are added to the spatial geometric encoding to obtain the correlations between channels.
@@ IV-D Implementation Details of MVGT
In this section, we describe the overall architecture of the model, including spatial encodings and the Transformer encoder, as illustrated in Fig. 1. For better optimization, we first apply GraphNorm [44] to normalize the input features to a range between 0 and 1. Subsequently, we perform geometric and regional structure encodings to obtain multi-domain embeddings. The encodings could be characterized as below:
$\bm{X}^{\prime}_{s}$ $=\text{GraphNorm}(\bm{X}_{s})$ (10)
$\bm{H}^{0}_{s}$ $=\text{SpatialEncoding}+\text{Proj}(\bm{X}^{\prime}_{s})$ (11)
We employ a Pre-LN Transformer structure, applying layer normalization (LN) before the multi-head attention (MHA) and the FFN. A recent study suggests that the Pre-LN structure yields more stable gradients and is more favorable for optimizer, enabling faster convergence [45] compared to Post-LN. Additionally, we utilize dropout to mitigate overfitting. This process is represented as follows:
$\bm{H}^{\prime l}_{s}$ $=\text{MHA}(\text{LN}(\bm{H}^{l-1}_{s}))+\bm{H}^{l-1}_{s}$ (12)
$\bm{H}^{l}_{s}$ $=\text{FFN}(\text{LN}({\bm{H}^{\prime l}_{s}}))+\bm{H}^{\prime l}_{s}$ (13)
Inspired by [30, 46], we feed the outputs recursively into the same modules, denoted as "recycling" in Fig. 1. The iterative refinement progressively refines the model’s ability to discriminate encoded information and understand emotional patterns, thereby helping the model capture more effective details.
@@ V-A Datasets
For our experiments, we select the SEED [5] and SEED-IV [47] datasets to evaluate the effectiveness of our model. These datasets consist of EEG signals recorded from subjects watching emotion-eliciting videos.
SEED dataset comprises data from 15 subjects who participated in three sessions, each separated by at least one week. Each session consists of 15 trials capturing emotional labels, with the emotion labels being positive, negative, and neutral.
SEED-IV dataset is constituted by EEG signals from 15 subjects across three separate sessions conducted at different times, using the same device as the SEED dataset. This dataset encompasses four emotion labels: neutral, sad, fear, and happy. In each session, each subject underwent 24 trials.
@@ V-B Settings
To prevent potential data leakage that could arise from segment-wise shuffling, we split the training and test sets at the trial level. Following the settings of previous studies [5, 10, 12, 13, 14, 16, 47, 31], we use pre-computed differential entropy (DE) features for the recognition task. For the SEED dataset, we use the first 9 trials of each subject as the training set and the last 6 trials as the test set, as done in previous research. The DE features are computed using five frequency bands extracted from 1s nonoverlapping windows. The model performance is evaluated based on the average accuracy and standard deviation across all subjects over two sessions of EEG data. Similarly, for the SEED-IV dataset, we use the first 16 trials as the training set and the last 8 trials as the test set. The DE features for SEED-IV are calculated using 4s windows.
Table I: The classification accuracies (mean/std) on SEED and SEED-IV. MVGT-L, MVGT-G, MVGT-H, MVGT-F: MVGT using LOBE, GENERAL, HEMISPHERE and FRONTAL schemes.
@@
The performance of our model is assessed using data from all three sessions.
For the input data, we use overlapping sliding windows of size $T$ along the time axis to extract sample fragments, with $T$ being set to 5. During experiments, the hidden dimension is set to 64 and the number of Gaussian basis functions is 32. The number of MHA layers is 4 and the number of attention heads is 2. The iterative refinement process is performed three times. We set the batch size to 32 and the learning rate within the range of 3e-5 to 3e-3. Cross-entropy is used as the loss function, and AdamW [48] is employed as the optimizer with a weight decay rate of 0.1.
V-C Results Analysis
We compare the classification results based on the SEED and SEED-IV datasets with recent state-of-the-art models, as shown in Table I. It is evident that our proposed model significantly outperforms the baseline models under the same experimental settings. For the SEED dataset, the model adopting the FRONTAL scheme achieves the best performance, with a classification accuracy of 96.45%. The LOBE scheme also achieves a slightly superior accuracy of 95.36%, compared to other models. For the SEED-IV dataset, the classification accuracy under the GENERAL scheme is 93.57%, achieving the best performance compared to baseline models. The MVGT model also demonstrates strong performance under other division schemes. Overall, our model achieves the best classification accuracy compared to the baselines. The results also suggest that selecting the specific division scheme relevant to the emotion task could enhance the expressive power of MVGT.
Figure 3: Confusion matrices of MVGT. (a) Confusion matrix of MVGT-F on the SEED dataset. (b) Confusion matrix of MVGT-G on the SEED-IV dataset. Each row of the matrix represents the true labels while each column serves as the predicted labels.
@@ V-C Results Analysis
Fig. 3 illustrates the confusion matrices of MVGT-F on the SEED and MVGT-G on the SEED-IV, respectively. The values represent the classification accuracy of the model for different emotion classes. For the SEED dataset, our model achieves the highest accuracy in recognizing positive emotion (98.12%), followed by neutral emotion (96.38%), with negative emotion being slightly lower (94.73%). Only 0.33% of positive emotion samples are misclassified as negative, while only 0.76% of negative emotion samples are recognized as positive, indicating the model’s effectiveness in distinguishing valence changes. For the SEED-IV dataset, our model performs best in recognizing neutral emotion, with an accuracy of 95.90%, while its performance on happy emotion is slightly lower than the other three emotions, with an accuracy of 90.76%. This could be attributed to the GENERAL scheme setting, making the model more sensitive to the balanced emotion.
Our model achieves state-of-the-art performance on both SEED and SEED-IV, primarily due to our comprehensive consideration of frequency, temporal, and spatial geometric information, combined with prior knowledge from neuroscience.
Figure 4: The learned inter-channel relationships from the SEED by the MVGT-F and from the SEED-IV by the MVGT-G are illustrated. The figures show the results of the last iteration in the iterative refinement, highlighting the top 10 channel pairs with the highest weights after softmax (darker colors indicate higher weights). Channels of the same brain region are represented in the same color. Rows correspond to attention heads, while columns represent the layers of the MHA mechanism.
@@ V-D Ablation Study
To validate the effectiveness of spatial encodings, we conduct ablation experiments on the SEED and SEED-IV datasets, as presented in Table II. By removing both types of spatial encodings, we repeat the aforementioned experiments under the same experimental settings. On the SEED dataset, the model achieves an accuracy of 93.79% with a standard deviation of 7.15%. Compared to MVGT-F, the accuracy decreases by 2.66% and the standard deviation increases by 2.75% after removing spatial encodings. For the SEED-IV dataset, the accuracy drops by 4.08%, resulting in 89.49%, with the standard deviation rising by 1.80% to 10.40%, when compared to MVGT-G. The experiments demonstrate that incorporating spatial structure information benefits the model performance in emotion recognition tasks. Under experimental settings that consider only geometric structure or brain region structure, the model’s classification accuracy improves over the plain model without any spatial encoding. Evidently, when considering both types of spatial structures simultaneously, the model performance significantly outperforms that of the plain model and models only using single spatial information. This indicates the effectiveness of our proposed spatial encodings and confirms that the expressive power of the graph transformer relies on the structural and positional encodings.
@@ V-E Visualization of Inter-channel relations
To better illustrate the correlations between channels, we visualize the inter-channel relations of MVGT-F on the SEED and MVGT-G on the SEED-IV. Given that the inter-channel relations might vary among different subjects, we calculate the average weights across all subjects. We focus on the last iteration of iterative refinement and select the 10 strongest connections of channel pairs. Fig. 4 shows the visualization results, where the rows represent the attention heads and the columns represent the layers of the MHA.
The parameters based on the SEED dataset indicate that emotion patterns are reflected in the activities of multiple brain regions. In the first layer of MVGT-F, the channels in the left frontal region have higher participation in the first attention head, while the channels in the right frontal region are more involved in the second head, potentially corresponding to positive and negative emotion patterns [9], respectively. In the second layer of the model, the parietal and occipital regions show higher involvement, which aligns with the findings on emotion patterns in [49]. As the model depth increases, the symmetrical connections in the lateral temporal regions of both hemispheres are enhanced, consistent with previous research by [5, 13, 16]. For the SEED-IV dataset, the connections in the frontal, parietal, and occipital regions are the most active, consistent with the findings of [13]. In the first attention head of MVGT-G, the strongest correlation is between O1 and PO3, followed by P4 and P2. Other connections are mainly distributed in the temporal and frontal regions. In the second head, the channel pairs (O1, PO5), (CB1, PO7), and (PO5, PO7) contribute the most to emotion recognition. Additionally, the connection between AF3 and FP1 provides important information for emotion processing, which aligns with the conclusions of [13, 16].
Overall, our model does not focus solely on the local information of a single brain region but instead considers both intra-regional and inter-regional information in depth. This confirms that emotional states result from interactions among widely distributed functional networks in the brain, as discussed by [50].
@@ VI Conclusions
In this paper, we propose a multi-view graph transformer based on spatial relations for EEG-based emotion recognition. Our model integrates information from multiple perspectives, including temporal, frequency and spatial domains. We incorporate spatial geometric encoding and brain region encoding to enhance the graph transformer’s ability to perceive spatial structures. Additionally, the model adaptively learns inter-channel relationships through the attention mechanism and the encoding of channel geometry. Extensive experiments on public emotion recognition datasets demonstrate that our proposed model outperforms other competitive baseline models.
Furthermore, the analysis of channel correlations indicates that emotional activities in the brain are not confined to a single local region but result from the coordinated action of multiple brain areas. Information from frontal, parietal, occipital, and lateral temporal lobes is valuable for emotion recognition to varing extents.